2 particles of equal mass m at A and B connected by a rodAB lying on smooth horizontal table.An impulse J applied at A in plane of table and perpendicular to AB.Find velocity of particle A.

Grade:12th pass

1 Answers

Sayan Bhowmik
6 years ago
As we know the impulse is change in momentum so, J=p2-p1=2(mv)sin(angle). Here angle is 90°. So J=4mv. V=J/(4m)